OpenStudy (ttop0816):

What is the equation of the line that passes through (1, 2) and is parallel to the line whose equation is 2x + y - 1 = 0? 2x - y - 4 = 0 2x + y - 4 = 0 << i chose this 2x + y + 4 = 0

OpenStudy (anonymous):

looks good again

OpenStudy (ttop0816):

how about this one?? What is the equation of the line with an x-intercept of -2 and a y-intercept of 1? y = -1/2x + 1 y = 1/2x + 1 << y = 2x + 1